1+1=3 2+2=8 3+3=15 4+4=24 5+6=?? CHECK YOUR GENIOUS

1+1=3
2+2=8
3+3=15
4+4=24
5+6=??
CHECK YOUR GENIOUS

Grade:11

3 Answers

imran
6 years ago
Hi CHOWDAIAH
 I Think answer is 36
      Because by observing above lines , i can say that
                 2*1+1=3
                 3*2+2=8
                 4*3+3=15
                 5*4+4=24
     then     6*5+6=36
             OK THANK Q
Lakshmipathi
6 years ago
HAII Choudaiah the correct answer is 36.........................................................................................................
YASWANTH
6 years ago
HI CHOWDAIAH THE ANSWER IS 47 ..I.E, 1+1*2=3: 2+2*3=8 …....5+6*7=47  TRY TO ANSWER AND UPLOAAD THESE TYPE OF QUESTIONS TOOOO